the integral of ∫(1/lnx)dx is.......its 1 divided by natural log of x w.r.t. dx

Grade:

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear kshitij

you can find the solution of these type of integral in terms of infinite series

I=∫(1/lnx)dx

let lnx =t

     or x= et

        dx =etdt

I =∫(1/t) etdt

  =∫(1/t)[1+t+t2/2! + t3/3! +..........]dt

  =∫[1/t +1+t/2! + t2/3! +..........]dt

  =lnt + t + t2/2*2! + t3/3* 3! +.........

  =ln(lnx) + lnx + (lnx)2/2*2! + (lnx)3/3* 3! +.........

  =ln(lnx) +  ∑(lnx)r/r*r!  where r varies from 1 to infinity

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

ankit singh
3 years ago

et lnx =t

     or x= et

        dx =etdt

I =∫(1/t) etdt

  =∫(1/t)[1+t+t2/2! + t3/3! +..........]dt

  =∫[1/t +1+t/2! + t2/3! +..........]dt

  =lnt + t + t2/2*2! + t3/3* 3! +.........

  =ln(lnx) + lnx + (lnx)2/2*2! + (lnx)3/3* 3! +.........

  =ln(lnx) +  ∑(lnx)r/r*r!  where r varies from 1 to infinity